# A Strategy for Preparing Quantum Squeezed States Using Reinforcement Learning

### v2

@@ Abstract
We propose a scheme leveraging reinforcement learning to engineer control fields for generating non-classical states. It is exemplified by the application to prepare spin-squeezed states for an open collective spin model where a linear control field is designed to govern the dynamics. The reinforcement learning agent determines the temporal sequence of control pulses, commencing from a coherent spin state in an environment characterized by dissipation and dephasing. Compared to the constant control scenario, this approach provides various control sequences maintaining collective spin squeezing and entanglement. It is observed that denser application of the control pulses enhances the performance of the outcomes. However, there is a minor enhancement in the performance by adding control actions. The proposed strategy demonstrates increased effectiveness for larger systems. Thermal excitations of the reservoir are detrimental to the control outcomes. Feasible experiments are suggested to implement the control proposal. The extension to continuous control problems and another quantum system are discussed. The replaceability of the reinforcement learning module is also emphasized. This research paves the way for its application in manipulating other quantum systems.
@@ I Introduction
Machine-learning techniques are emerging as effective tools in physics Murphy2012 ; Sutton2018 ; Nature549195 , and among them, reinforcement learning (RL) offers the potential to optimize the control field for high-dimensional, multistage processes in complex scenarios. Deep reinforcement learning (DRL) can provide a control strategy to engineer the dynamics as long as the evolution follows certain differential equations. In physics, many optimal problems can be treated as control problems of finding means to steer a system to achieve a certain target. The search for optimal control fields can be formulated as RL tasks Sutton2018 ; MURPHY2018 ; PRX8031084 . A semiquantum reinforcement learning approach is employed to adapt a qubit to an unknown state by successive single-shot measurements and feedback AQT21800074 . A four-step strategy combining constant-value and time-varying controls results in an enhanced longtime stable spin squeezing for a collective spin system coupled to a bosonic field PRA103032601 .
@@
This paper is organized as follows: In Sec.II, we delineate the reinforcement-learning-based framework to prepare nonclassical states. In Sec.III, the RL module employed within the control scheme is shown. In Sec.IV, we explicate the quantum model for the generation of spin-squeezed states. In Sec.V, we present the procedure to prepare squeezed states of the open collective spin system via reinforcement learning. In Sec.VI, we check the performance of this method including the influence of the frequency of applying the control pulses, the granularity of control actions, the scalability across various sizes, and the impact of thermal excitation. In Sec.VII, we discuss the feasible experiments to implement this control scheme based on BEC, the application to continuous control situations and the extension to another quantum system. The replaceability of the RL agent is also discussed. Finally, we conclude in Sec.VIII.
@@ IV Collective Spin Model
where $\hat{J}_{\vec{n}_{i}}$=$(\hat{J}_{x},\hat{J}_{y},\hat{J}_{z})\cdot\vec{n}_{i}$, ($i=1,2$) and $\vec{n}_{1}=(-\sin{\phi},\cos{\phi},0)$, $\vec{n}_{2}=(\cos{\theta}\cos{\phi},\cos{\theta}\sin{\phi},-\sin{\theta})$. The collective spin operator $\hat{J}_{s}$=$(\hat{J}_{x},\hat{J}_{y},\hat{J}_{z})\cdot(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, where $\theta$ = $\arccos(\frac{\langle J_{z}\rangle}{|{\hat{J}}|})$, $\phi$ = $\mathrm{sign}(\langle\hat{J}_{y}\rangle)\arccos(\frac{\langle\hat{J}_{x}\rangle}{|{\hat{J}}|\sin\theta})$, the norm $|{\hat{J}}|=\sqrt{\langle\hat{J}_{x}\rangle^{2}+\langle\hat{J}_{y}\rangle^{2}+\langle\hat{J}_{z}\rangle^{2}}$ PR50989 ; RMP90035005 ; JPB39559 . Here the direction $\vec{n}_{\perp}$=$\vec{n}_{1}\cos\varphi+\vec{n}_{2}\sin\varphi$ with
@@
where $A\equiv\langle J_{\vec{n}_{1}}^{2}-J_{\vec{n}_{2}}^{2}\rangle,\text{ \ }\;B\equiv\langle\left[\hat{J}_{\vec{n}_{1}},\hat{J}_{\vec{n}_{2}}\right]_{+}\rangle$.
@@
as the squeezing parameter. Here $\Delta\hat{J}_{z}^{2}$= $\langle\hat{J}_{z}^{2}\rangle-\langle\hat{J}_{z}\rangle^{2}$, indicates the spin squeezing in the z-direction. For the RL, we use the reverse of $\xi_{Z}^{2}$ to set the reward-punishment rule during the control process. The reward function we use is $R=10\langle\Delta R_{t}\leq 0\rangle-\langle\Delta R_{t}<0\rangle$, where
=$\langle\bullet\rangle=\left\{\begin{array}[]{ll}-1&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ \bullet\leq 0,\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ 1&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ \bullet>0,\end{array}\right.$ (7)
$\Delta R_{t}=\frac{1}{\xi_{z}^{2}}_{(t+1)}-\frac{1}{\xi_{z}^{2}}_{(t)}$. $\frac{1}{\xi_{z}^{2}}_{(t)}$ is the inverse of the squeezing parameter at the sample time $t$ during training. This reward function means that the agent gets 10 points if the action makes the squeezing parameter decrease, whereas deduct 1 point if the squeezing parameter rises. Subsequently, in accordance with the fundamental principles of RL, the agent will steer the direction with minimum squeezing parameter approaching the $z$ direction, namely, $\varphi=\pi/2$ in Eq. 5. Compared to using the reverse of $\xi_{R}^{2}$ with variable $\theta$ and $\phi$, the choice of $\xi_{Z}^{2}$ possesses a more direct physical interpretation since $\hat{J}_{z}$ signifies the population imbalance between the two modes within the Ramsey interferometer mentioned above PRA46R6797 ; PRA475138 ; RMP90035005 .
@@ V.1 Initial coherent spin state
Figure 2: (a) Evolution of the spin squeezing parameter (6) versus training epochs while 100 segments of actions are evenly taken across each evolution time interval $[0,2]$. Every 30 training epochs are shown within 600 consecutive training rounds. The subfigure in (a1) shows the evolution of the corresponding averaged QFI defined as $\bar{F}$=$[F(\rho,\hat{J}_{x})+F(\rho,\hat{J}_{y})+F(\rho,\hat{J}_{z})]/(3N^{2})$. (b) Top view of the 3-action ($\Omega(t)=2,0,-2$) square control pulse corresponding to the spin squeezing curves in (a). (c) 30 samples (differentiated by $n$) of the squeezing parameter at $t=2$ of those in (a). As shown in Figure (a), there are control sequences causing $\xi_{Z}^{2}$ to diverge at the beginning. However, such control sequences are invalid. Therefore, we express instances where $\xi_{Z}^{2}>5$ uniformly as 5 to facilitate the visualization of the results in $(c)$. (d) The squeezing parameters versus time for the constant coherent control ($\Omega(t)=-2$) and the RL-designed control, with the corresponding degree of mixing indicated by $Tr[\rho^{2}(t)]$. The subscript $RL$ means $\Omega(t)$ obtained by reinforcement learning, $Con$ means $\Omega(t)=-2$.(e1) evolution of the angle $\varphi$ in Eq. (5).(e2) The control fields $\Omega(t)$ designed by the RL agent corresponding to $(d)$ and the constant control field $\Omega(t)=-2$.(f1)and(f2) are the Husimi representation of the initial CSS and the spin squeezed state at $t=2$.(f3) and (f4) are the polar plots of the Wigner functions for the states. We use $N=2J=20$ hereafter. The decaying parameters are $\gamma=0.001\kappa$.
@@ V.2 Prepare spin-squeezed states by machine-designed pulses
VI Control Result
@@
We conjecture that the application frequency of the pulses impacts the outcomes of the control. To investigate the evolution of the squeezing parameter versus different application frequencies, we discretize the temporal interval [0, 2] into different numbers of segments, the more segments, the more frequently applying the pulses. It can be seen in Fig. 3(a), with increasing the segments, the squeezing parameter is depressed more stationary and lower. Even more, with increasing of the control pulses, the variance of the squeezing parameter is also depressed more obviously. Straightforwardly, there is a contradiction between the control performance and operation difficulty. One may conjecture the number of control actions also influences the control performance. However, as shown in Fig. 3(b), there is no obvious advantage for more control actions with the same maximum control amplitude in this control.
@@ VIII Conclusion
In conclusion, we have proposed a reinforcement-learning-based control strategy to generate non-classical collective spin states in an open environment. The machine learning agent can design a suite of control sequences to prepare spin-squeezed states accompanied by entanglement. We find larger frequency of applying the pulses enhances the performance of the control strategy. However, more choices of actions do not contribute obviously to the control outcomes. The scalability of this framework is demonstrated, facilitating its applicability to larger systems. Notably, thermal excitations progressively undermine the control performance. The control proposal can be realized within current experiments of atomic Bose-Einstein condensates. The extension of the control paradigm to continuous control problems and other quantum systems is discussed. The versatility of this scheme is underlined by the potential of substituting the reinforcement learning agent with other optimization modules.

### v4

@@ Abstract
We propose a scheme leveraging reinforcement learning to engineer control fields for generating non-classical states. It is exemplified by the application to prepare spin-squeezed states for an open collective spin model where a linear control field is designed to govern the dynamics. The reinforcement learning agent determines the temporal sequence of control pulses, commencing from a coherent spin state in an environment characterized by dissipation and dephasing. Compared to the constant control scenario, this approach provides various control sequences maintaining collective spin squeezing and entanglement. It is observed that denser application of the control pulses enhances the performance of the outcomes. However, there is a minor enhancement in the performance by adding control actions. The proposed strategy demonstrates increased effectiveness for larger systems. Thermal excitations of the reservoir are detrimental to the control outcomes. Feasible experiments are suggested to implement this control proposal based on the comparison with the others. The extensions to continuous control problems and another quantum system are discussed. The replaceability of the reinforcement learning module is also emphasized. This research paves the way for its application in manipulating other quantum systems.
@@ I Introduction
Machine-learning techniques are emerging as effective tools in physics Murphy2012 ; Sutton2018 ; Nature549195 , and among them, reinforcement learning (RL) offers the potential to optimize the control field for high-dimensional, multistage processes in complex scenarios. Deep reinforcement learning (DRL) can provide a control strategy to engineer the dynamics as long as the evolution follows certain differential equations. In physics, many optimal problems can be treated as control problems of finding means to steer a system to achieve a certain target. The search for optimal control fields can be formulated as RL tasks Sutton2018 ; MURPHY2018 ; PRX8031084 . A semiquantum reinforcement learning approach is employed to adapt a qubit to an unknown state by successive single-shot measurements and feedback AQT21800074 . A four-step strategy combining constant-value and time-varying controls results in an enhanced longtime stable spin squeezing for a collective spin system coupled to a bosonic field PRA103032601 . A measurement-based adaptation protocol is introduced to optimize a quantum state for maximum overlap with an unknown reference state using quantum reinforcement learning PRA98042315 . This study provides a foundational framework for the development of quantum reinforcement learning protocols leveraging quantum data.
@@
This paper is organized as follows: In Sec.II, we delineate the reinforcement-learning-based framework to prepare nonclassical states. In Sec.III, the RL module employed within the control scheme is shown. In Sec.IV, we explicate the quantum model for the generation of spin-squeezed states. In Sec.V, we present the procedure to prepare squeezed states of the open collective spin system via reinforcement learning. In Sec.VI, we check the performance of this method including the influence of the frequency of applying the control pulses, the granularity of control actions, the scalability across various sizes, the impact of thermal excitation, and the comparison with other current protocols. In Sec.VII, we discuss the feasibility of this control scheme in terms of BEC experiments, the application to continuous control situations and the extension to another quantum system. The replaceability of the RL agent is also discussed. Finally, we conclude in Sec.VIII.
@@ IV Collective Spin Model
where $\hat{J}_{\vec{n}_{i}}$ = $(\hat{J}_{x},\hat{J}_{y},\hat{J}_{z})\cdot\vec{n}_{i}$, ($i=1,2$) and $\vec{n}_{1}=(-\sin{\phi},\cos{\phi},0)$, $\vec{n}_{2}=(\cos{\theta}\cos{\phi},\cos{\theta}\sin{\phi},-\sin{\theta})$. The collective spin operator $\hat{J}_{s}$ = $(\hat{J}_{x},\hat{J}_{y},\hat{J}_{z})\cdot(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$, where $\theta$ = $\arccos(\frac{\langle\hat{J}_{z}\rangle}{|{\hat{J}}|})$, $\phi$ = $\mathrm{sign}(\langle\hat{J}_{y}\rangle)\arccos(\frac{\langle\hat{J}_{x}\rangle}{|{\hat{J}}|\sin\theta})$, the norm $|{\hat{J}}|=\sqrt{\langle\hat{J}_{x}\rangle^{2}+\langle\hat{J}_{y}\rangle^{2}+\langle\hat{J}_{z}\rangle^{2}}$ PR50989 ; RMP90035005 ; JPB39559 . Here the direction $\vec{n}_{\perp}$ = $\vec{n}_{1}\cos\varphi+\vec{n}_{2}\sin\varphi$ with
@@
where $A\equiv\langle\hat{J}_{\vec{n}_{1}}^{2}-\hat{J}_{\vec{n}_{2}}^{2}\rangle,\text{ \ }\;B\equiv\langle\left[\hat{J}_{\vec{n}_{1}},\hat{J}_{\vec{n}_{2}}\right]_{+}\rangle$.
@@
as the squeezing parameter. Here $\Delta\hat{J}_{z}^{2}$ = $\langle\hat{J}_{z}^{2}\rangle-\langle\hat{J}_{z}\rangle^{2}$, indicates the spin squeezing in the z-direction. For the RL, we use the reverse of $\xi_{Z}^{2}$ to set the reward-punishment rule during the control process. The reward function we use is $R=10\langle\Delta R_{t}\geq 0\rangle-\langle\Delta R_{t}<0\rangle$, where
={ $\langle\bullet\rangle=\left\{\begin{array}[]{ll}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ 0&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ \bullet False,\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ 1&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ \bullet Truth,\end{array}\right.$ (7)
$\Delta R_{t}=\frac{1}{\xi_{z}^{2}}_{(t+1)}-\frac{1}{\xi_{z}^{2}}_{(t)}$. Here $\frac{1}{\xi_{z}^{2}}_{(t)}$ is the inverse of the squeezing parameter at the sample time $t$ during training. This reward function means that the agent gets 10 points if the action makes the squeezing parameter decrease, whereas deduct 1 point if the squeezing parameter rises. Subsequently, in accordance with the fundamental principles of RL, the agent will steer the direction with minimum squeezing parameter approaching the $z$ direction, namely, $\varphi=\pi/2$ in Eq. (5). Compared to using the reverse of $\xi_{R}^{2}$ with variable $\theta$ and $\phi$, the choice of $\xi_{Z}^{2}$ possesses a more direct physical interpretation since $\hat{J}_{z}$ signifies the population imbalance between the two modes within the Ramsey interferometer mentioned above PRA46R6797 ; PRA475138 ; RMP90035005 .
@@ V.1 Initial coherent spin state
Figure 2: (a) Evolution of the spin squeezing parameter (6) versus training epochs while 100 segments of actions are evenly taken across each evolution time interval $[0,2]$. Every 30 training epochs are shown within 600 consecutive training rounds. The subfigure in (a1) shows the evolution of the corresponding averaged QFI defined as $\bar{F}$ = $[F(\rho,\hat{J}_{x})+F(\rho,\hat{J}_{y})+F(\rho,\hat{J}_{z})]/(3N^{2})$. (b) Top view of the 3-action ($\Omega(t)=2,0,-2$) square control pulse corresponding to the spin squeezing curves in (a). (c) 30 samples (differentiated by $n$) of the squeezing parameter at $t=2$ of those in (a). As shown in Fig. (a), there are control sequences causing $\xi_{Z}^{2}$ to diverge at the beginning. However, such control sequences are invalid. Therefore, we express instances where $\xi_{Z}^{2}>5$ uniformly as 5 to facilitate the visualization of the results in (c). (d) The squeezing parameters versus time for the constant coherent control ($\Omega(t)=-2$) and the RL-designed control, with the corresponding degree of mixing indicated by $Tr[\rho^{2}(t)]$. The subscript RL means $\Omega(t)$ obtained by reinforcement learning, Con means $\Omega(t)=-2$. (e1) evolution of the angle $\varphi$ in Eq. (5). (e2) The control fields $\Omega(t)$ designed by the RL agent corresponding to (d) and the constant control field $\Omega(t)=-2$. (f1)and(f2) are the Husimi representation of the initial CSS and the spin squeezed state at $t=2$. (f3) and (f4) are the polar plots of the Wigner functions for the states. We use $N=2J=20$ hereafter. The decaying parameters are $\gamma=\gamma_{z}=0.001\kappa$.
@@ V.2 Prepare spin-squeezed states by machine-designed pulses
VI Control Results
VI.1 The results of the control method
@@
We conjecture that the application frequency of the pulses impacts the outcomes of the control. To investigate the evolution of the squeezing parameter versus different application frequencies, we split the temporal interval [0, 2] into different numbers of segments, the more segments, the more frequently applying the pulses. It can be seen in Fig. 3(a), with increasing the segments, the squeezing parameter is depressed more stationary and lower. Even more, with increasing of the control pulses, the variance of the squeezing parameter is also depressed more obviously. Straightforwardly, there is a contradiction between the control performance and operation difficulty. One may conjecture the number of control actions also influences the control performance. However, as shown in Fig. 3(b), there is no obvious advantage for more control actions with the same maximum control amplitude in this control.
@@
VI.2 The comparison with current protocols
In the two-component weakly interacting Bose-Einstein condensate Nature40963 , the minimum value for the squeezing parameter $\xi^{2}=N(\Delta\hat{J}_{\vec{n}_{1}})^{2}/(\langle\hat{J}_{\vec{n}_{2}}\rangle^{2}+\langle\hat{J}_{\vec{n}_{3}}\rangle^{2})$ approaches $\frac{1}{2}{\left(3/N\right)}^{2/3}$ at an optimized direction. In the work Nature4551216 , a BEC of thousands 87Rb atoms can be described by the one-axis twisting Hamiltonian $\hat{H}=E_{C}/2\hat{J}_{z}^{2}-2E_{J}/N\hat{J}_{x}$, where the Josephson energy $E_{J}$ characterizes the tunneling between the two condensates, and the charging energy $E_{C}$ describes the repulsive interaction inside BEC. The squeezing parameter $\xi^{2}_{N}=\xi_{R}^{2}\cos\varphi^{2}=\Delta\hat{J}_{z}^{2}/\Delta\hat{J}_{z,{\rm ref}}^{2}$ is used to describe squeezing, where $\varphi$ is the phase difference of the two modes. The concurrent presence of number squeezing and high phase coherence enabled the spin squeezing parameter $\xi_{R}^{2}=-3.8$ dB for the two main well pairs of a six-well lattice (approximately 2,200 atoms), and $\xi_{R}^{2}=-2.3$ dB for the double-well configuration (approximately 1,600 atoms). In a similar BEC of 87Rb (total number 2,300) nonlinear Ramsey interferometer experiments Nature4641165 , the minimal number squeezing factor $\xi_{N}^{2}=4\left(\Delta\hat{J}_{z}\right)^{2}/N=-8.2_{-1.2}^{+0.8}\,\text{dB}$ after taking the technical noise due to coupling-pulse imperfections and magnetic field fluctuations into account. In another work Nature4641170 , spin squeezing was generated in an atom chip by controlling the elastic collision interactions, resulting to $\xi_{R}^{2}=-2.5_{0.6}^{-0.6}\,\text{dB}$ for 1250$\pm$45 atoms.
Theoretically, a scheme using repeated Rabi pulses transforming one-axis-twisting model to two-axis-twisting type with the Hamiltonian $\hat{H}(t)=\chi\hat{J}_{z}^{2}+\Omega(t)\hat{J}_{y}$, and the squeezing parameter $\xi^{2}=2(\Delta\hat{J}_{min})^{2}/J=4(\Delta\hat{J}_{min})^{2}/N$ can approach the Heisenberg limited $1/N$ PRL107013601 . In another protocol using the chopped random basis technique PRA93013851 , the squeezing parameter $\xi^{2}=N(\Delta\hat{J}_{x})^{2}/|\langle\hat{J}_{z}\rangle|^{2}\sim 2.1/N^{0.94}$ with a time-random-fluctuation Hamiltonian. In a four-step strategy including constant-value and time-varying controls to prepare spin squeezing for an ensemble of the two-level systems coupled to a single-mode bosonic field, when $N=2J=8$, the squeezing parameter $\xi^{2}=2(\Delta\hat{J}_{min})^{2}/J$ can approach $-8$ dB PRA103032601 . In these proposals, the optimized spin squeezing occurs at certain time and phase angles.
According to our results in Fig. 3, Fig. 4, and Fig. 5, the squeezing parameter $\xi_{Z}^{2}=N\Delta\hat{J}_{z}^{2}/|\langle\hat{J}_{s}\rangle|^{2}$ can be stabilized in the range of $-5$ dB to $-7$ dB for a long time. Compared with the current works, the advantage of our scheme lies in its ability to provide spin-squeezed states along $\hat{J}_{z}$ over a longer time window without the need to extract squeezed states at a specific squeezing time. Moreover, we have considered quantum mechanical dissipation, making our theoretical research closer to actual experimental scenarios.
@@ VIII Conclusion
In conclusion, we have proposed a reinforcement-learning-based control strategy to generate non-classical collective spin states in an open environment. The machine learning agent can design a suite of control sequences to prepare spin-squeezed states accompanied by entanglement. We find larger frequency of applying the pulses enhances the performance of the control strategy. However, more choices of actions do not contribute obviously to the control outcomes. The scalability of this framework is demonstrated, facilitating its applicability to larger systems. Notably, thermal excitations progressively undermine the control performance. The control proposal can be realized within current experiments of atomic Bose-Einstein condensates. The extensions of the control paradigm to continuous control problems and other quantum systems are discussed. The versatility of this scheme is underlined by the potential of substituting the reinforcement learning agent with other optimization modules.